Sam is fencing his yard that is 42 feet long and 120 feet wide. He wants to put fencing around the entire perimeter of his yard. The fencing costs $4.25 per foot. How much will it cost to fence his yard?
Answer by ikleyn(52786)

.

Calculate the perimeter  P = 2*(42+120) = 324 feet.


Then multiply it by the price per feet to get the answer $4.25*324 = 1377 dollars.